find f(1/4) where
f(x) = 12 sin{pi x + 3pi/4}

Answer by lwsshak3(11628)

find f(1/4) where
f(x) = 12 sin{pi x + 3pi/4}
f(1/4)=12*sin(π/4+3π/4)
=12*sin(π)
=12*0
=0